### v1

# Turning Trash into Treasure: Accelerating Inference of Large Language Models with Token Recycling

Abstract
The rapid growth in the parameters of large language models (LLMs) has made inference latency a fundamental bottleneck, limiting broader application of LLMs. Speculative decoding represents a lossless approach to accelerate inference through a guess-and-verify paradigm, leveraging the parallel capabilities of modern hardware. Some speculative decoding methods rely on additional structures to guess draft tokens, such as small models or parameter-efficient architectures , which need extra training before use. Alternatively, retrieval-based train-free techniques build libraries from pre-existing corpora or by n-gram generation. However, they face challenges like large storage requirements, time-consuming retrieval, and limited adaptability. Observing that candidate tokens generated during the decoding process are likely to reoccur in future sequences, we propose Token Recycling. This approach stores candidate tokens in an adjacency matrix and employs a breadth-first search (BFS)-like algorithm on the matrix to construct a draft tree. The tree is then validated through tree attention. New candidate tokens from the decoding process are then used to update the matrix. Token Recycling requires <2MB of additional storage and achieves approximately 2x speedup across all sizes of LLMs. It significantly outperforms existing train-free methods by 30% and even a training method by 25%. It can be directly applied to any existing LLMs and tasks without the need for adaptation.
Figure 1: A comparison of typical speculative decoding and Token Recycling. Typical speculative decoding drafts some tokens and verifies them in parallel in one decoding step. Unlike speculative decoding which discards candidate tokens, Token Recycling stores them in an adjacency matrix. In future generations, draft tokens are retrieved from the matrix which is updated with new candidate tokens. Token Recycling effectively recycles tokens in the decoding process.
1 Introduction
Large Language Models (LLMs) (Brown et al. 2020; Gemini Team et al. 2023; Touvron et al. 2023; Meta 2024) have made significant contributions to the development of artificial intelligence, becoming the foundation of numerous applications such as chatbots, code assistants, agents (OpenAI 2023; Chen et al. 2021; Wang et al. 2024a). However, due to the auto-regressive decoding strategy, LLMs can only generate a single token at each decoding step, leading to slow response times for queries. The latency mainly comes from transferring billions of parameters from High-Bandwidth Memory to the accelerator’s cache at each decoding step, rather than arithmetic computations (Kim et al. 2024; Shazeer 2019; Cai et al. 2024).
Many approaches (Xu et al. 2024; Frantar and Alistarh 2023; Dao 2024; DeepSeek-AI 2024) seek to reduce the latency, with speculative decoding as a key lossless technique. This approach employ a guess and verify process to obtain multiple tokens during a single decoding step  (Chen et al. 2023; Leviathan, Kalman, and Matias 2023; Miao et al. 2024; Xia et al. 2023). It first speculates several subsequent draft tokens and then verifies them using the original LLMs. The time cost of verification on multiple tokens is comparable to that of generating one token due to the high parallel performance of accelerators. Once some draft tokens are correct, the whole decoding steps can be greatly reduced while maintaining the quality. To fully utilize the parallel capability of accelerators, tree attention slightly adjust the attention mask to verify multiple draft tokens continuations in one model forward (Cai et al. 2024; Miao et al. 2024).
Speculative decoding aims not only to maintain quality but also to minimize the cost of speculation. Some researchers construct additional model structures to guess the draft tokens, including small draft models (Leviathan, Kalman, and Matias 2023; Chen et al. 2023) and parameter-efficient structures (Cai et al. 2024; Lin et al. 2024). However, these approaches require resources for additional training on each LLM. The typical approach to achieve train-free speculative decoding is retrieve-based. In this case, a retrieval library is pre-defined to obtain tokens following the suffix of current content as draft tokens. Several methods have been proposed in this category, each with its trade-offs: (i) REST (He et al. 2023) transforms existing corpora into a retrieval library, but the storage is large, retrieval is time-consuming, and the library lacks flexibility as it’s static to any queries. (ii) PLD (Saxena 2023) only retrieves the previous content with minimal cost. However, it can not predict new tokens or new combinations of tokens. (iii) Lookhead (Fu et al. 2024) construct and update an n-gram library through LLMs decoding n times. However, LLMs have to generate n-grams while in inference, causing low efficiency.
Furthermore, all these speculative decoding approaches fail to fully utilize candidate tokens, which are multiple possible next tokens generated by LLMs at each decoding step. Typically, only the top-1 candidate token from accepted tokens is selected as the output, while other candidate tokens, including all candidate tokens from rejected tokens, are discarded, such as ‘k’ and ‘keys’ in Figure 1. However, we observe that when current input tokens reappear in future generations, the following tokens could be candidate tokens generated several steps prior. Based on the observation, we propose Token Recycling, a method which utilizes candidate tokens as draft tokens. It stores candidate tokens in an adjacency matrix. Before each decoding step, a BFS-like approach retrieves a draft tree from the matrix, which is then verified using tree attention. Once verified, the newly generated candidate tokens update the matrix. (i) The matrix provides a flexible retrieval library that is tailored to each query and offers low retrieval costs due to its small size (<2MB). (ii) Compared to using the previous content solely, candidate tokens naturally include more tokens, providing many possible continuations. (iii) The construction and update of our library (matrix) utilize the ‘trash’ tokens without requiring any additional generation.
We conduct comprehensive experiments on general benchmark SpecBench (Xia et al. 2024), and specialized dataset on code domain, MBPP  (Chen et al. 2021) with different sizes of Vicuna (Zheng et al. 2023) and Code Llama (Roziere et al. 2023) (7b, 13b, 33b/34b). The results show that Token Recycling greatly exceeds previous train-free approaches, and improves more than 31% on all sizes. The speed-up ratio even exceeds the widely used training approach–Medusa, demonstrating its high efficiency.
Our contributions are summarized below:
  • We observe that candidate tokens can be reused as draft tokens in subsequent sequences. Based on this observation, we propose a train-free speculative decoding method, Token Recycling, which saves candidate tokens and retrieves draft tokens based on the adjacency matrix.
  • The matrix requires minimal storage space (<2MB) with a low retrieval cost, and covers many new tokens. Continuously updating provides a dynamic retrieval space, but doesn’t need additional generation.
  • The experimental results show that Token Recycling achieves approximately 2x speedup on all sizes of LLMs. It improves more than 31% compared to previous train-free approaches and achieves a new SOTA, even exceeding a training approach.
2 Background
In this section, we overview the speculative decoding. We begin with a formal definition of auto-regressive decoding. Next, we explore speculative decoding, highlighting the guess-and-verify strategy. Finally, we introduce tree attention, which verifies multiple draft sequences simultaneously.
2.1 Auto-Regressive Decoding
Auto-Regressive (AR) is the default decoding strategy of LLMs. At each decoding step $t$, LLMs calculate the probability distribution of the next token given the current content $s=(x_{0},x_{1},\cdots,x_{t})$ which $x_{i}\in\mathcal{V}$:
$p_{t+1}=P(x|s;\theta).$
Here $\mathcal{V}$ represents the vocabulary and $\theta$ represents the parameters of the LLM. The next token is selected from $p_{t+1}$ depending on the sampling methods. Followed Kou et al. (2024), we focus on greedy decoding in this paper. The next token is as follows:
$x_{t+1}=\text{argmax}\ p_{t+1}.$
Candidate tokens refer to the top-$k$ tokens with the highest probabilities
$(x_{t+1}^{0},x_{t+1}^{1},\ldots,x_{t+1}^{k-1})=\text{argtop}k(p_{t+1})$
where $k$ is the number of candidate tokens. Here, $\text{argtop}k(\cdot)$ is a custom operator that returns the indices of the top-$k$ highest values in the probability distribution $p_{t+1}$.
Figure 2: An overview of Token Recycling. The adjacency matrix, initialized by inheriting from the previous query, stores candidate tokens. Token Recycling first retrieves a draft tree from the matrix based on the last token of the current content. The tree is then compressed into a merged sequence with a corresponding tree attention mask and sent to the LLM for a forward pass. After processing, all possible draft sequences are extracted and verified. The longest correct sequence is selected and added to the content. Finally, the top-k candidate tokens are used to update the matrix for the next iteration.
2.2 Speculative Decoding
Through the guess and verify approach, speculative decoding effectively utilizes the parallel capability of GPUs to accelerate inference. Given $s$, it first guesses $n$ subsequent draft tokens $(\tilde{x}_{t+1},\cdots,\tilde{x}_{t+n})$. The combination $(s,\tilde{x}_{t+1},\cdots,\tilde{x}_{t+n})$ is then sent to LLMs for one forward pass, resulting in:
$p_{t+1}$ $=P(x\mid s;\theta),$
$\tilde{p}_{t+i}$ $=P(x\mid s,\tilde{x}_{t+1},\ldots,\tilde{x}_{t+i-1};\theta),\quad\text{for }i=2,\ldots,n.$
$p_{t+1}$ is the same as AR decoding so the ground-truth $x_{t+1}$ can be selected using ‘argmax’. If the draft token $\tilde{x}_{t+1}$ matches $x_{t+1}$, then $\tilde{p}_{t+2}$ is assumed to identical to $p_{t+2}$ in AR decoding. Thus, the next token is selected by: $x_{t+2}=\text{argmax}\ \tilde{p}_{t+2}$. This verify process continues until the draft token does not match the selected token, indicated by:
$x_{t+j}=\text{argmax}\ \tilde{p}_{t+j}\neq\tilde{x}_{t+j}.$
Ultimately, $j$ new tokens are confirmed in one forward pass. The time cost of one forward pass with $(s,\tilde{x}_{t+1},\cdots,\tilde{x}_{t+n})$ is nearly the same as with $s$ due to the high parallel performance of accelerators. Figure 1 shows an example. Consider the current content ending with the token ‘for’ and the draft tokens are [‘i’, ‘in’, ‘range’, ‘(’]. After the forward pass, the output token for ‘for’ is ‘i’, and for ‘i’ is ‘in’, but for ‘in’ is ‘zip’, which fails to match the guessed ‘range’. ‘i’ and ‘in’ are accepted tokens while the others are rejected. Despite ‘zip’ not matching ‘range’, it is still a valid output generated by the LLM based on the correct content, resulting in three tokens [‘i’, ‘in’, ‘zip’] being confirmed in one forward pass. Overall, speculative decoding effectively accelerates inference while maintaining the same output as AR decoding.
2.3 Tree Attention
Traditional causal attention masks are designed for linear sequences, where each token attends to all previous tokens. This design restricts speculative decoding to verifying one sequence at a time. However, as the sequence lengthens during draft token generation, the number of potential continuations increases. For example, in the draft tree in Figure 2, the token following ‘guest’ could be ‘speaker’ or ‘speak’, while both ‘at’ and ‘for’ could follow ‘speaker’. This creates a need to verify multiple draft sequences simultaneously.
Tree attention modifies the attention mask to address this. It compresses multiple sequences into a single merged sequence, such as [‘guest’, ‘speaker’, ‘speak’, ‘at’, ‘for’, ‘ings’], while preserving a tree structure in the tree attention mask. Each child node attends only to its parent nodes, preventing sibling tokens from interfering with each other. After the LLM processes the merged sequence, all possible sequences such as ‘guest speaker’, ‘guest speaker at’, ‘guest speaker for’, and ‘guest speak’ and so on, along with their corresponding output tokens are extracted based on the tree structure and verified in parallel. The longest correct sequence is selected as the final output. Tree attention enables speculative decoding to verify multiple draft sequences in one decoding step, significantly enhancing efficiency while maintaining output quality.
In summary, speculative decoding, through guess and verify and tree attention, improves the inference latency robustly and efficiently compared to AR decoding.
3 Methodology
Figure 2 provides an overview of Token Recycling. It leverages a hot-start adjacency matrix to store candidate tokens and employs a BFS-like algorithm to construct a draft tree. It utilizes tree attention to verify draft sequences and continuously updates the matrix with new candidate tokens generated during the decoding process.
3.1 Adjacency Matrix Initialization
The adjacency matrix $\mathcal{M}$ is a key component in Token Recycling, used to store top-$k$ candidate tokens for each token in the vocabulary:
$\mathcal{M}\in\mathcal{V}^{|\mathcal{V}|\times k}$
where $k$ is a user-defined hyperparameter. Each element $\mathcal{M}[i,j]$ indicates that the token $V_{M[i,j]}$ is the $j$-th candidate token associated with $V_{i}$. The use of matrix format, as opposed to other structures like tries, enables efficient parallel processing of candidate tokens, which is crucial for reducing retrieval and update times.
Initially, all elements are set to zero, meaning that a token must appear in draft tokens before it has valid candidate tokens. This initialization leads to the matrix starting with limited predictive capability, potentially causing inefficiencies during the early stages of inference. To mitigate this limitation, we implement a hot start strategy. This involves inheriting the matrix from the previous query, thereby leveraging existing knowledge. Even if queries differ in the domain, candidate tokens often include common expressions and patterns that frequently appear across various queries. Nevertheless, as the first query lacks a preceding query, the matrix is set to zero. Consequently, hot start ensures that the matrix has a broader starting point, covering a wide range of potential continuations.
3.2 Draft Tree Retrieval
The adjacency matrix $\mathcal{M}$ stores candidate tokens, which can be used as draft tokens when their corresponding tokens appear later. Directly using the matrix could only determine the immediate next token, such as finding ‘speaker’ following ‘guest’ (see Figure 2). Even if ‘speaker’ is correct, it only slightly improves upon AR decoding, adding just one additional token. In fact, the matrix also holds possible continuations for these candidate tokens, suggesting subsequent tokens like ‘at’ following ‘speaker’. Extending the sequence step by step allows for longer draft sequences. Furthermore, by storing top-$k$ candidate tokens, multiple potential continuations can be explored in parallel for each token, such as ‘at’, and ‘for’ following ‘speaker’. This breadth-first search (BFS)-like process enables the exploration of the sequence length and diverse continuations while constructing a draft tree with the adjacency matrix. Inspired by Cai et al. (2024), the tree is static and imbalanced. The complete tree structure is shown in Appendix 8.1.
  • Static: The number of children for each node remains constant across all decoding steps. This fixed structure facilitates pre-processing, such as preparing attention masks, and enables efficient parallel operations during layer traversal. By avoiding the need to individually traverse each node, it significantly reduces retrieval time.
  • Imbalance: Nodes positioned earlier in each layer have more children and extend deeper. Since candidate tokens are ordered by their probabilities in the matrix, this design prioritizes the most probable continuations, allocating computational resources to higher confident paths.
The BFS-like approach for retrieving the draft tree begins with the matrix $\mathcal{M}$ and the tree structure $Tree$. The retrieval process starts from the root, which is the last token of current content. Take the token ‘guest’ in Figure 2 as an example. As the root forms the first layer, all candidate tokens for ‘guest’ are extracted from $\mathcal{M}$, resulting in [‘speaker’, ‘speak’, ‘Spe’]. According to $Tree$, the first layer allows each token to have two children, Therefore, ‘speaker’ and ‘speak’, which have the top-2 probabilities, are selected and added to the second layer. The process then proceeds to expand a new layer starting from [‘speaker’] and [‘speak’]. All their candidate tokens are retrieved in parallel, resulting in [‘at’, ‘for’, ‘is’] and [‘ings’, ‘in’, ‘ers’]. $Tree$ specifies that the first node in this layer (‘speaker’) can have two children, while the subsequent node (‘speak’) can only have one child. Consequently, the new layer tokens are [‘at’, ‘for’], and [‘ings’]. This process repeats until the specified depth is reached. Finally, a merged sequence $S$ is constructed through traversing the draft tree by layers. The detailed Algorithm 1 is provided in Appendix 8.1.
This retrieval method achieves the desired length and variety of the draft sequences, resulting in a single merged sequence that can be sent to the LLM.
3.3 Verification and Update
Once the merged sequence $S$ is constructed, it is combined with the pre-processed attention mask into the LLM for a forward pass. The verification aligns with Section 2.3. All potential draft sequences are extracted based on $Tree$, and the longest correct sequence is selected as the final output.
Following verification, the adjacency matrix $\mathcal{M}$ is updated in parallel based on the output distributions $\tilde{p}_{i+1}$ of each draft token $x_{i}\in S$:
$\mathcal{M}[\tilde{x}_{i}]=\text{argtop}k(\tilde{p}_{i+1}).$
It is important to note that duplicates may be in $S$ since multiple preceding tokens may have the same candidate token. Consequently, the output distributions of duplicates are likely to differ. When performing updates in parallel, CUDA operations may merge these updates, leading to variations in the final result. For example, if $x_{i}$ appears twice and has two different top-2 output tokens, $[y_{0},y_{1}],[z_{0},z_{1}]$, then $\mathcal{M}[x_{i}]$ could be updated to exactly one of the following results: $[y_{0},z_{1}]$, $[y_{0},y_{1}]$, $[z_{0},z_{1}]$ or $[z_{0},y_{1}]$. We do not resolve this merging, as adding controls reduces overall performance, as discussed later in Section 5.2.
The update process directly overwrites the previous candidate tokens, and leverages the new ones as draft tokens for subsequent decoding steps. This allows the retrieval space to dynamically adapt to the current content, focusing on the most relevant and probable continuations. It also eliminates the necessity for extra operations beyond the standard decoding to update the retrieval space.
#Para Method SpecBench MBPP
MT Trans Sum QA Math RAG MAT Ts/s Speed MAT Ts/s Speed
7B Vanilla 1.00 1.00 1.00 1.00 1.00 1.00 1.00 54.30 1.00 1.00 56.15 1.00
Lookahead 1.42 1.12 1.21 1.21 1.52 1.13 1.64 69.03 1.27 1.66 79.16 1.41
PLD 1.53 0.98 2.36 1.10 1.50 1.74 1.75 83.30 1.53 1.39 66.65 1.19
REST 1.37 1.05 1.12 1.42 1.06 1.30 1.84 66.29 1.22 2.08 87.08 1.55
Medusa* 1.90 1.57 1.48 1.58 1.87 1.45 2.31 89.41 1.65 - - -
Ours 2.17 1.90 1.94 1.95 2.40 1.78 2.70 110.06 2.03 2.93 131.20 2.34
shuffle 2.16 1.91 1.93 1.94 2.37 1.72 2.71 109.07 2.01 - - -
13B Vanilla 1.00 1.00 1.00 1.00 1.00 1.00 1.00 39.41 1.00 1.00 41.31 1.00
Lookahead 1.29 1.06 1.16 1.12 1.48 1.09 1.63 47.50 1.21 1.73 56.87 1.38
PLD 1.45 1.01 2.10 1.02 1.55 1.65 1.67 57.01 1.45 1.48 52.20 1.26
REST 1.51 1.14 1.31 1.50 1.17 1.50 1.82 53.34 1.35 2.05 70.13 1.70
Medusa* 1.94 1.66 1.57 1.62 1.98 1.53 2.39 67.92 1.72 - - -
Ours 1.98 1.77 1.89 1.75 2.21 1.73 2.72 74.57 1.89 3.08 93.42 2.26
33B Vanilla 1.00 1.00 1.00 1.00 1.00 1.00 1.00 18.44 1.00 1.00 19.44 1.00
Lookahead 1.32 1.09 1.20 1.17 1.55 1.14 1.61 23.03 1.25 1.70 29.22 1.50
PLD 1.43 1.06 1.94 1.08 1.55 1.41 1.55 25.89 1.40 1.41 25.89 1.33
REST 1.63 1.27 1.42 1.61 1.29 1.57 1.81 26.99 1.46 2.10 36.85 1.90
Medusa* 1.98 1.75 1.63 1.68 2.09 1.61 2.32 33.11 1.80 - - -
Ours 1.95 1.75 1.92 1.77 2.24 1.78 2.63 35.16 1.91 3.05 45.43 2.34
Table 1: Performance comparison of different methods on SpecBench (Vicuna) and on MBPP (Code Llama) across three parameter sizes. Vanilla represents AR decoding. Metrics include Mean Accepted Token (MAT), Tokens per Second (Ts/s), and Speedup ratio relative to Vanilla (Speed). Categories for SpecBench include cover Multi-turn Conversation (MT), Translation (Trans), Summarization (Sum), Question Answering (QA), Mathematical Reasoning (Math), and Retrieval-Augmented Generation (RAG), with Speed as the displayed metric. MBPP results exclude Medusa as it lacks a Code Llama variant. Medusa* involves training while others are training-free. shuffle indicates performance after shuffling the data order (Section 5.2).
In summary, Token Recycling capitalizes on the “trash” present in speculative decoding by implementing a cycling process between candidate and draft tokens. It accelerates inference without the need for additional model structures or training, making it highly adaptable and seamlessly integrated with any architecture or model size.
4 Experiment
4.1 Experimental Setup
Align with previous work  (Kou et al. 2024), we focus on common computational redundancy scenarios, specifically greedy decoding with a batch size of one. The following evaluation metrics are used: Mean Accepted Token (MAT) (Xia et al. 2024) represents the average number of tokens confirmed in a single decoding step; Tokens per Second (Ts/s) measures the number of tokens processed per second; Speedup ratio compares the performance relative to HuggingFace’s implementation of AR decoding. We set $k=8$ for $\mathcal{M}$ (<2MB storage in sum) and the draft tree structure is shown in Appendix 8.1. All experiments are conducted using Pytorch 2.3 with a single A100-80GB GPU and 128 CPUs under CUDA 12.2.
Datasets and LLMs
We conduct experiments on SpecBench (Xia et al. 2024) and MBPP (Austin et al. 2021). SpecBench is a comprehensive benchmark encompassing diverse scenarios including Multi-turn Conversation, Translation, Summarization, Question Answering, Mathematical Reasoning, and Retrieval-Augmented Generation. MBPP is a widely used dataset in code generation, a domain with a growing demand for efficient generation. These datasets enable a comparative analysis with prior work across both general and specialized domains. For the LLMs, we follow the standard practice of utilizing Vicuna (Chiang et al. 2023) for SpecBench and Code Llama (Roziere et al. 2023) for MBPP across three different scales: 7B, 13B, and 33B [^1] The largest model of Code Llama is 34B, for consistency and convenience in our comparisons, we refer to it as 33B..
Baseline
We compare Token Recycling with three lossless, train-free, and retrieval-based methods. Lookahead constructs an n-gram retrieval library through additional n-gram generation during decoding, consuming significant computational resources. PLD treats previous content as the retrieval library, which is constrained and cannot introduce new tokens or new token combinations. REST builds the retrieval library from existing training datasets, requiring large storage and considerable retrieval time. The static nature of the library also prevents it from adapting to individual queries. Furthermore, we include a parameter-efficient tuning baseline for border comparison. Medusa adds multiple additional LM heads in the final layer to predict draft tokens. We focus on losses Medusa-1 since Medusa-2 is lossy. All baselines use their default hyperparameters.
4.2 Main Results
Figure 3: Effects of tree breadth and depth on MAT and Tokens/s in (a) and (b), and the impact of updating strategies in (c).
Table 1 shows the performance of Token Recycling compared to other methods. On SpecBench, it achieves more than a 2x speedup on the 7B model, nearly 30% higher than the previous train-free methods. Even compared to tuning Medusa, it shows an improvement of almost 25%. For the 13B and 33B models, it consistently provides nearly 2x speedup, maintaining the 30% acceleration advantage. These results demonstrate that Token Recycling is the most effective train-free method on SpecBench, offering substantial and consistent speedup across all model sizes.
Notably, Token Recycling achieves the best speedup across most sub-tasks as well, except it slightly trails PLD on Summarization. This may be due to this task often involves many repetitions of previous content. However, the performance gap between Token Recycling and PLD narrows as the model size increases, reaching only a 1% difference with the 33B model. This is due to larger models tending to generate new tokens rather than repeat previous content. In other tasks such as Multi-turn Conversation, Translation, Question Answering, and Mathematical Reasoning, Token Recycling shows a significant improvement of about 40%~70% for the 7B model. This demonstrates the strong generalization of our method across various scenarios. Although the improvement on Retrieval-Augmented Generation is less than 3% for the 7B model, it increases with model size, exceeding 10% for the 33B one. This improvement is consistent with the preference of larger models for new tokens. Compared to the general domain, all methods achieve greater acceleration on the code domain due to its higher content redundancy. Token Recycling provides approximately 2.3x speedup across all model scales, achieving the SOTA performance.
Furthermore, performances on Trans show the advantages of our method compared to PLD and REST. While PLD shows negligible speedup (close to 1x) and REST achieves its lowest speedup across tasks, Token Recycling consistently delivers over 1.75x speedup across all model sizes. Notably, on the 7B model, PLD results in a slowdown, and REST achieves just 1.05x, whereas Token Recycling reaches 1.9x. Trans requires generating new tokens continuously, involving minimal repetition of previous content. Additionally, it is highly context-sensitive, making it challenging to find exact matches from any pre-existing database. These pose challenges for PLD and REST. In contrast, the adaptive and diverse retrieval space of Token Recycling leads to superior performance. In addition to Speed, Token Recycling achieves the highest MAT across both benchmarks. This is attributed to its shorter retrieval times and the avoidance of additional generations like Lookahead. This allows for deeper and wider draft trees, enabling more tokens to be accepted in a single decoding step.
Token Recycling demonstrates significant improvements across all scenarios, highlighting its efficiency and broad applicability. Importantly, our method is train-free and self-drafting, allowing for an approximate 2x speedup that can be seamlessly applied as a “free lunch” to any existing LLM.
5 Analysis
5.1 Tree Structure
As previously outlined in Section 3.2, our tree structure is static and imbalanced. The tree size is a crucial factor to accelerate. A larger tree allows more tokens confirmed in one decoding step but also introduces more computational overhead, increasing the time required for each decoding step. To investigate the impact of tree size, specifically its depth and breadth on acceleration performance, experiments are conducted on MT-Bench using Vicuna-7B.
Breadth
Increasing the breadth of the tree allows for covering more possibilities during decoding. In Figure 3(a), the breadth is expanded by adding nodes while keeping the depth fixed at six layers. This leads to a consistent improvement in MAT. However, when the breadth exceeds 80, Tokens/s begins to decrease. The additional computational overhead eventually outweighs the benefits of a higher MAT.
Depth
Increasing the depth of the tree allows for accepting longer sequences during decoding. In Figure 3(b), with the number of nodes fixed at 80, the depth is gradually increased. MAT initially rises rapidly but eventually shows minimal improvement, while Tokens/s noticeably fluctuates. Because the matrix stores candidate tokens for only adjacent steps, longer sequences weaken the connections between distant tokens. This limitation reduces the effectiveness of increased depth, causing Tokens/s to fluctuate.
5.2 Ablation Study
Data Order
In Token Recycling, the adjacency matrix is initialized with hot start, meaning it inherits the matrix from the previous query. By default, SpecBench organizes data so that instances from the same domain are processed consecutively. This sequential ordering might allow domain-specific patterns to persist, potentially optimizing token retrieval and boosting performance. To evaluate the robustness of hot start, we conduct an experiment where all data instances are shuffled across different domains. This shuffling ensures that each query is independent, removing any potential domain influence from the previous query. As shown in Table 1 under shuffle, the performance impact is minimal, with the Speed decreasing by less than 1%. This demonstrates that hot start is effective and robust, even without domain continuity between queries.
Update Strategies
Section 3.1 discuss duplicate tokens in the merged sequence during matrix updates. We compare three updating strategies: using candidate tokens from the first occurrence, from the last occurrence, and the current method (merging via parallel CUDA operations). Figure 3(c) indicates that using the last occurrence yields the highest MAT, which may benefit from more contextual information. However, the differences among different strategies in MAT are minimal. In terms of Tokens/s, the current approach significantly outperforms the other two, as it avoids the additional processing required to manage token positions, thereby reducing delays. Speculative decoding is highly sensitive to latency, any extra operation must provide substantial benefits to outweigh its time cost.
Figure 4: Time allocation for each operation when LLMs response to a query.
5.3 Time Allocation
For speculative decoding to be effective, it is essential to maintain a high hit rate while minimizing the time spent on additional operations. We divide each decoding step into several components: preprocessing , retrieving draft tokens, model forward pass, verifying draft sequences, and updating the matrix, input tokens, and key-value cache. The average time spent on each component is shown in Figure 4. The results indicate that the majority of the time is consumed by the model forward pass. The verification process also takes a significant amount of time due to the need to extract and verify all feasible paths. Retrieving draft tokens and updating operations take roughly the same amount of time.
6 Related Work
Efficient inference is crucial for real-time applications and low-resource scenarios. Many strategies have been developed to reduce latency (Zhou et al. 2024b). Among these, speculative decoding (Chen et al. 2023; Leviathan, Kalman, and Matias 2023; Miao et al. 2024; Xia et al. 2023) has emerged as a losses technique for achieving lossless speedup by predicting multiple possible continuations simultaneously, thus reducing the number of decoding steps needed without compromising accuracy. Some speculative decoding methods rely on additional draft models to guess draft tokens. These typically involve using smaller models from the same series (Zhao et al. 2024; Spector and Re ; Sun et al. 2023; Liu et al. 2024b; Yuan et al. 2024; Gong et al. 2024) or training new models with a shared vocabulary (Leviathan, Kalman, and Matias 2023; Chen et al. 2023; Zhou et al. 2024a; Li et al. 2024). Additionally, Kou et al. (2024); Wang et al. (2024b) propose training the original LLMs to enable non-aggressive decoding. While effective, these approaches require managing or training multiple models, which can be non-trivial and resource-intensive. Other methods focus on parameter-efficient structures. These approaches minimize the need for complete retraining but still require model-specific training and adaptation, limiting their scalability and general applicability (Lin et al. 2024; Liu et al. 2024a).
There are also train-free methods that construct retrieval libraries to obtain draft tokens (Yang et al. 2023). Lookahead (Fu et al. 2024) uses the LLM to generate n-grams through multiple decodings, building a retrieval library that can hit multiple tokens in one step. Despite this advantage, it requires the LLM to generate n-grams while responding to queries, which reduces efficiency. PLD (Saxena 2023) retrieves only from previous content, resulting in minimal overhead and significant speedup in high-redundancy tasks like summarization. However, it provides little acceleration for tasks requiring the generation of new content, like translation. REST (He et al. 2023) constructs retrieval libraries using existing corpora and performs well in common scenarios. However, this approach requires large storage, time-consuming retrieval, and cannot adapt to each query.
Token Recycling is a train-free, retrieval-based method that stores candidate tokens in an adjacency matrix and reuses them as draft tokens. It requires no additional generation, covers a broader range of possible continuations, and demands minimal storage with low retrieval costs. The update process ensures the retrieval space remains adaptable
7 Conclusion
In this work, we introduce Token Recycling, a speculative decoding method for accelerating the inference of LLMs. It utilizes an adjacency matrix to store candidate tokens and retrieve a draft tree, which is then verified with tree attention. The matrix is updated with new candidate tokens generated during decoding. Token Recycling could be integrated seamlessly with existing LLMs and various tasks. As a train-free approach, it achieves a speedup of approximately 2x with <2MB of additional storage, improving over 31% compared to previous train-free approaches.